Status Report

NASA Space Station On-Orbit Status 23 September 2004

By SpaceRef Editor
September 23, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 23 September 2004

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2004) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.  

Update on Elektron:  The Elektron remains off at this time and troubleshooting continues.   [Moscow suspects that the high O2 pressure that shut down the machine yesterday again resulted from of the system reacting to an increase in hydrogen (H2) pressure (via the BRD pressure equalization unit).  The crew verified with a nitrogen purge that the 4mm-diameter H2 line is indeed congested, leading TsUP to conclude that the filter at the inlet to the solenoid valve just upstream of the vacuum vent to space is clogged with contaminants, probably crystallized alkali (KOH).  Today’s troubleshooting consists of attempts to clean the filter by disconnecting the connector upstream of the vacuum valve, filling the lines with water and, after reconnection and pressure checks of the line, allowing a vacuum venting to flush the lines and any associated residue.  During activation tests over the weekend, in order to avoid the need for repressurizing the O2 and H2 lines with N2 at this time, the machine will not be allowed to run for the length of time (11 hrs.) that would require this repress.]

Service Module Atmosphere Revitalization Subsystem, Book 2, Mission Operations Directorate, 9 October 2000 [Acrobat] According to this document’s introduction “This book contains information for the crew about procedures and rules for the atmosphere revitalization subsystem, Elektron, Vozdukh, Micropurification Unit, and Fire Detection and Suppression Subsystem operations, as well as their schematic and operation logic.”

To keep ppO2 (oxygen partial pressure) in the cabin atmosphere within its prescribed limits, another repressurization with Progress O2 is being prepared for next weekend.

The altitude reboost of the station was successfully conducted this morning at 8:05am EDT as scheduled, lasting 10 min and yielding a delta-V of 2.7 m/s.  This put the ISS in an ideal flight condition for the upcoming Soyuz launch and landing events.

Early in the morning, right after breakfast, FE/SO Michael Fincke deployed two acoustic dosimeters, one on each crewmember, for being worn for 24 hours (with a microphone on the shirt collar).  A third dosimeter was deployed in the Lab for a 24-hr. static data take (last time done: 8/5).   [Tonight, after about 15 hours of measurements, dosimeter data will be recorded and the hardware power-cycled, for another data take tomorrow morning after 8.5-hr. sleep.  At that point, the crew will deploy the dosimeters statically in the station for the duration of the day, record measurements tomorrow night and stow the instruments.  Acoustic data must be taken twice per Increment, each time for the duration of the 16-hour crew workday.]

Padalka was busy most of the day setting up and then periodically monitoring the transfer of the stored potable water from the Progress 15P cargo ship to the Service Module (SM) Rodnik tankage.   [First Gennady hooked up hoses and the GZhS air/liquid separator unit for transferring the water while removing air bubbles, then initiated the automated transfer of the water, first from Progress tank #1 (BV1) to SM tank #1, then from one #2 Rodnik tank to the other.  Each of the two BV tanks required 5 hrs of pumping, and Gennady monitored the GZhS for air bubbles every hour for the first 4 hrs and every 20 min. during the last pumping hour, periodically relieving air from the GZhS as required.  Late in the evening, he is to terminate the activity and tear down the transfer hardware.]

Initiating a new round of the monthly preventive maintenance of Russian segment air ventilation systems, Mike Fincke worked one hour in the Funktsionalnyi-Grusovoi Blok (FGB) for uninstalling and replacing the FGB’s two dust collector filters (PS1 & PS2).   [Last time done: 8/13.]

CDR Padalka completed the periodic servicing task of changing out ASU toilet system equipment, then checked out the ASU.   [Replaced with new units were the U-receptacle (MP) and filter insert (F-V).  The old items were stowed for disposal.]

Gennady transferred the accumulated data files of his BIO-5 Rasteniya-2 (“Plants-2”) plant growth experiment to the Packet laptop for subsequent downlink to the ground via Regul comm.  Topping off its water container as required was also one of his tasks today.

Later, the CDR conducted an inventory/audit of a range of utensils (SPP) stowed in SM and FGB locations, to support planning for future deliveries.

Mike Fincke continued the changeout of service life-expired CMOS (complementary metal-oxide semiconductor) batteries of onboard laptops, today on the SSC Router (shell #6007), another IBM ThinkPad 760XD that was originally scheduled for R&R on 9/1.

The FE also successfully transferred the triaxial accelerometer unit of the IWIS (internal wireless instrumentation system) from its temporary installation near the TVIS chasses behind SM panel 137, back to the Lab, mounting it in the module’s forward endcone.  As part of the task, he also relocated RSU (remote sensing unit) #1027 from the Node to the Lab and installed RSU #1026 (from stowage) in the Node instead.   [The accelerometer was in the SM in support of TVIS troubleshooting, now complete.]

Afterwards, Fincke completed his final session for the general MedOps PFE (periodic fitness evaluation) that checks up on blood pressure and electrocardiogram during programmed exercise on the CEVIS cycle ergometer in the Lab.  Readings were taken with the BP/ECG (blood pressure/electrocardiograph) and the HRM (heart rate monitor) on a chest strap.  Padalka assisted his crewmate as CMO (crew medical officer) with electrode placement, etc.   [BP/ECG and HRM provide automated noninvasive systolic and diastolic blood pressure measurements while also monitoring and displaying accurate heart rates on a continual basis at rest and during exercise.]

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

Later, Gennady continued periodic checkout work on the Russian TVS television system, today on the Klest (KL-103Ts) TV camera.   [The Klest was turned on and off by the SM’s automated daily timeline system (SPP), and Gennady monitored its test imagery on the #2 monitor (VKU2).]

In support of a two-day NASA/JSC simulation of 12A/Increment 11 being conducted by Shuttle (9/22) and ISS (today), the Lab RWS (robotics workstation) of the SSMRS (Space Station Remote Manipulator System) was powered up by the ground at 9:00am and kept active until 1:00pm.

For the EarthKAM (Earth Knowledge Acquired by Middle school students, EK) payload, Mike performed maintenance on its DCS 460 digital camera by recharging the batteries in the camera to prolong their life and prepare them for the next operational session in late October.

Fincke did the routine maintenance of the SM’s SOZh environment control & life support system and Padalka prepared the regular IMS (inventory management system) delta file for export/import to the IMS databases.

Soyuz Crew: ‘Phone Home’

“Soyuz crews have been provided with a Iridium/Motorola-9505 satellite phone and a Garmin GPSMAP 76 handheld GPS unit. Both units have the ability to function anywhere on Earth.”

The CDR retrieved the Motorola-9505 Iridium satellite phone from its location in the Soyuz TMA-4 descent module (DM) for its monthly recharging of the lithium-ion battery and started the 30-min. process.   [For safety, before powering up the recharge unit (BV-5), the telephone, as before (8/19/04), was placed in a single CTB (crew transfer bag), which then was contained in a triple CTB.  The crew inspected the two CTBs to ensure their integrity (internal damage to CTB zippers tested for Iridium battery recharging on the ground had allowed an intentionally triggered fire to penetrate through the zippers).  The charging was monitored without taking the satphone out of the containment.  Upon completion, Padalka removed the phone, placed it inside its SSSP Iridium kit and stowed it back in the DM’s operational data files (ODF) container.]

The crew completed their regular daily physical exercise program on TVIS (now cleared for use without the 160-lb load restriction, but still speed-constrained until a future IFM), RED exerciser and VELO cycle with load trainer.

The crew’s regular (once every two weeks) teleconference with the JSC Astronaut Office (Kent Rominger), via S-band S/G, was deferred.

At 5:15am, Mike Fincke conducted a ham radio session with students at Investigator Science and Technology Centre in Adelaide, Australia, via “telebridge” over Hawaii.   [The Centre is a South Australian not-for-profit center committed to creating enthusiasm for science, technology and engineering among young people.  The guest school for this contact was Kilburn Primary School, with an enrolment of approximately 130 students (~20% Aboriginal, ~25% from non-English speaking backgrounds).  Questions by the students, aged 10-11 and in classes 5 & 6, were uplinked beforehand.]

For his next “Saturday Science” program (9/25), the Science Officer has now selected another ISSI (In-Space Soldering Investigation) test.

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

Yesterday, the Loop A TRRJ (Thermal Radiator Rotary Joint) test was completed successfully. Test activities included the String 2 Blind Mode with PPL (pre-positioned load) targets, switchover from String 2 to String 1, and moding to Shutdown.  Later, after this morning’s reboost, testing resumed on Loop B with TRRJ software initialization and mode to Autotrack.

The extended testing of the Russian ASN-M satellite navigation system, which will run through 9/25 and requires daily telemetry monitoring, is continuing, controlled by the SM’s automated timeline sequencer (SPP).   [When functioning, the ASN will use GLONASS satellites (the Russian GPS equivalent) to update the state vector (SV, position & velocity data) without using the ground (which up to now has to uplink daily SV updates) or requiring SV transfers from the U.S. segment (USOS) from time to time.]

For the Uragan earth-imaging program listed today on his discretionary task list, Padalka again had a number of opportunities for using the Kodak 760 DSC (digital still camera) with 800mm-lens from SM windows #9, now available again in LVLH attitude, on selected targets of interest.   [Today’s targets were Lake Lop Nor (to answer the question whether it actually exists or not), the cities of Samarkand and Tashkent, the Chervak water reservoir and river valley, the eruption of the volcano Etna in Sicily, incursion of Atlantic Ocean water into the Mediterranean via Gibraltar, Madrid, and the mountain kingdom of Andorra.]

Expedition 9 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

Today’s CEO photo targets, in the current LVLH attitude no longer limited by restrictions on the use of the Lab nadir/science window, were Amazon River Mouth, Atlantic Ocean (weather over the Atlantic NE of the Amazon River was predicted to be clear for internal wave photography. The sunglint point was almost directly below the station), Navassa Island reef, Caribbean (high resolution photography of this reef is necessary to assess its current morphology, extent, and ecological health.  The reef was located to the left of track), and Hurricane Ivan Damage (Dynamic Event.  This overpass provided opportunities to photograph damage from Hurricane Ivan along the eastern Appalachian Mountains.  Such photography will be useful for characterizing the post-flooding hydrologic dynamics of rivers, lakes, and ephemeral standing water bodies).

CEO images can be viewed at these websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 9 crew visit:

Current Soyuz 9S (Expedition 10 + 1) flight plan (Leroy Chiao, Salizhan Sharipov, Yuri Shargin):

  • Launch — 10/11, 12:17am ET (Moscow: 8:17am; Baikonur: 10:17am)
  • Docking @ DC1 — 10/13, 2:05am ET (Moscow: 10:05am)
  • Hatch Opening (docking + 2 orbits) — 10/13, 5:05am EDT (Moscow: 1:05pm).

Current 8S (Expedition 9 + 1) flight plan (Gennady Padalka, Michael Fincke, Yuri Shargin):

  • Hatch Closing — 10/19, 3:15pm ET (Moscow: 11:15pm; Kustanai: 10/20, 1:15am)
  • Undocking from FGB — 10/19, 6:20pm ET (Moscow: 10/20, 2:20am; Kustanai: 10/20, 4:20am)
  • Deorbit Burn — 10/19, 8:52pm ET (Moscow: 10/20, 4:52am; Kustanai: 10/20, 6:52am)
  • Landing — 10/19, 9:45pm ET (Moscow: 10/20, 5:45am; Kustanai: 10/20, 7:45am.

Other upcoming events:

  • Soyuz 9S relocate to FGB nadir port — 11/18;
  • Progress 15P undock – 11/23;
  • Progress 16P launch — 11/24;
  • EVA-12 — 12/28;
  • Progress 16P undock — 1/29/05;
  • Progress 17P launch — 1/30/05;
  • EVA-13 — 2/21/05;
  • Shuttle/LF1 launch — NET 3/6/05;
  • Shuttle/LF1 undock — NET 3/16/05.

 U.S. & Russian Segment Status  (as of today, 1:30pm EDT)
Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is Off.  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is On.  SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring.  MCA (major constituents analyzer) is Off.  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is Off.  SKV-2 is On (SM panel mods completed 4/8; SKV-2 activation failed 4/20; new replaceable condensate removal line installed on 9/9).  SFOG slot #2 fan suspect (not usable).
  • SM Working Compartment:  Pressure (mmHg) — 739; temperature (deg C) — 25.7; ppO2 (mmHg) — 149.8; ppCO2 (mmHg) — 2.9.
  • SM Transfer Compartment:  Pressure (mmHg) — 741; temperature (deg C) — 19.1.
  • FGB Cabin:  Pressure (mmHg) — 744; temperature (deg C) — 20.7.
  • Node:  Pressure (mmHg) — 736.2; temperature (deg C) — 22.6 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 738.4; temperature (deg C) — 23.2; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 738.5; temperature (deg C) — 21.8; shell heater temp (deg C) — n/a, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • (n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in directed position (dual angle/blind mode, non solar-tracking, biased for drag reduction).
  • SM batteries:  All batteries (8) are on line in “Partial Charge” mode.  
  • FGB batteries:  Battery #3 is off line; all other batteries (5) are on line in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

Command & Data Handling Systems (C&DH)

  • C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
  • GNC-1 MDM (vers. R4) is prime; GNC-2 (vers. R4) is backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off (backup).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

Propulsion System (PS):

  • Total propellant load available: 4158 kg (9167 lb) as of 9/16/04;  [SM(552) + FGB(3326) + Progress M(280)].  (Capabilities: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems (ACS):

  • 3 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2s RPC-17 failed 4/21/04; was replaced 6/30/04).
  • State vector source — RS
  • Attitude source — RS
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH XVV (local vertical/local horizontal = earth-fixed: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -7.2 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems (C&T):

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.